# Data-Adaptive Graph Framelets with Generalized Vanishing Moments for Graph Signal Processing

### v2

@@ II-A Preliminaries
$f=\sum_{i\in\mathcal{I}}\langle f,\phi_{i}\rangle_{\mathcal{H}}\phi_{i},\quad\forall f\in\mathcal{H}.$ (1)
@@
Proof.
It is obvious that (a)$\Rightarrow$(b). For (b)$\Rightarrow$(a), in view of $\bm{X}\bm{X}^{\top}=\bm{I}$ and $\bm{\Phi}\bm{\Phi}^{\top}=\bm{I}$, we have $\bm{A}\bm{A}^{\top}=\bm{I}$. The condition $\bm{B}\bm{A}^{\top}=\bm{0}$ follows from $\mathrm{span}(\bm{\Psi})\perp\mathrm{span}(\bm{\Phi})$. By the isomorphism between $\mathbb{R}^{m}$ and $\mathrm{span}(\bm{X})=\mathrm{span}(\bm{\Phi})\oplus\mathrm{span}(\bm{\Psi})$, we deduce that $\mathbb{R}^{m}=\mathrm{span}(\bm{A})\oplus\mathrm{span}(\bm{B})$, which implies the full rank condition $\mathrm{rank}(B)=m-m_{1}$ of $\bm{B}$. Hence (b)$\Rightarrow$(a). Consequently, (a)$\Leftrightarrow$(b).
Now suppose (a) holds. We next prove that (i) – (iv) are equivalent. In fact, the equivalence between (i) and (ii) is obvious. From [41, Lemma 1], we see that $\bm{B}\bm{B}^{\top}\bm{B}=\bm{B}$ is equivalent to $\{\bm{B}_{i,:}\}_{i\in[m_{2}]}$ is a tight frame of $\mathrm{span}(\bm{B})$, which is equivalent to $\{\bm{\psi}_{i}\}_{i\in[m_{2}]}$ is a tight frame of $\mathrm{span}(\bm{\Psi})$ in view of (b). Hence, (iii) and (iv) are equivalent. We only need to show (i) and (iii) are equivalent. Obviously, (i)$\Rightarrow$(iii) by the condition $\bm{B}\bm{A}^{\top}=\bm{0}$. Now given (iii), we see that $\bm{B}(\bm{A}^{\top}\bm{A}+\bm{B}^{\top}\bm{B})=\bm{B}$, which implies $\bm{B}(\bm{A}^{\top}\bm{A}+\bm{B}^{\top}\bm{B}-\bm{I})=\bm{0}$. By the full rank condition of $\bm{B}$, we deduce (i). Therefore, (i) – (iv) are equivalent. ∎
Lemma 1 can be extended to the following more general result, which will be useful in our later construction of graph framelet systems.
Corollary 1.
@@
$\mathbb{V}_{\bm{X}}$ $:=\mathrm{span}(\bm{X})\,\,with\,\,\bm{X}:=\llbracket\bm{X}_{i}\rrbracket_{i\in[m]},$ (2)
$\mathbb{V}$ $:=\mathrm{span}(\bm{\Phi})\,\,\,with\,\,\bm{\Phi}:=\llbracket\bm{\Phi}_{\ell}\rrbracket_{\ell\in[r]},$ (3)
$\mathbb{W}$ $:=\mathrm{span}(\bm{\Psi})\,\,\,with\,\,\bm{\Psi}:=\llbracket\bm{\Psi}_{\ell}\rrbracket_{\ell\in[r]}.$ (4)
Proof.
@@
$\left\{\begin{matrix}\mathcal{A}:=\mathrm{diag}(\bm{A},\ldots,\bm{A})\bm{E}\in\mathbb{R}^{m_{1}r\times mr},\\ \\ \mathcal{B}:=\mathrm{diag}(\bm{B},\ldots,\bm{B})\bm{E}\in\mathbb{R}^{m_{2}r\times mr}.\end{matrix}\right.$ (6)
Then, we have $\mathcal{A}\mathcal{A}^{\top}=\bm{I}$, $\mathcal{B}\mathcal{A}^{\top}=\bm{0}$, and $\mathcal{B}^{\top}\mathcal{B}=\bm{I}-\mathcal{A}^{\top}\mathcal{A}$. Moreover, $\bm{\Phi}=\mathcal{A}\bm{X}$ and $\bm{\Psi}=\mathcal{B}\bm{X}$. The conclusion then directly follows from applying Lemma 1. ∎
Remark 1.
We call the matrices $\bm{A},\bm{B}$ in Corollary 1 as filters or filter matrices, and the matrices $\bm{X}_{i},i\in[m]$ as subspace matrices. We formulate Corollary 1 as an algorithm in Algorithm 1. We highlight that the key of Corollary 1 is that each subspace $\mathrm{span}(\bm{X}_{i})$ has the same dimension $r$ for all $i\in[m]$. Such a condition will play a crucial role in our later construction of tight graph framelet systems.
@@
2:$\bm{\Xi}_{\ell}:=\text{{CONCAT}}\left(\{\bm{\xi}_{i,\ell}\}_{i=1}^{m}\right),\,\ell\in[r]$.
3:$\bm{\Phi}:=\text{{CONCAT}}\left(\left\{\bm{A}\bm{\Xi}_{\ell}\right\}_{\ell=1}^{r}\right)$.
4:$\bm{\Psi}:=\text{{CONCAT}}\left(\left\{\bm{B}\bm{\Xi}_{\ell}\right\}_{\ell=1}^{r}\right)$.
@@
II-B Construction of $\mathcal{V}$-Framelet Systems
@@
$\mathbb{V}_{j}=\mathbb{V}_{j-1}\oplus\mathbb{W}_{j-1},\quad\mathbb{V}_{j-1}\perp\mathbb{W}_{j-1},\quad j\in[J],$ (7)
then by doing the decomposition $J-1$ times, we have
$\mathbb{V}_{J}=\mathbb{V}_{0}\oplus\mathbb{W}_{0}\oplus\mathbb{W}_{1}\oplus\dots\oplus\mathbb{W}_{J-1}$ (8)
@@
$\mathbb{V}_{j}:=\oplus_{k=1}^{n_{j}}\mathbb{V}_{j,k},\quad\mathbb{W}_{j}:=\oplus_{k=1}^{n_{j}}\mathbb{W}_{j,k},$ (9)
@@
$\begin{gathered}\oplus_{k^{\prime}\in\mathcal{C}_{j,k}}\mathbb{V}_{j+1,k^{\prime}}=\mathbb{V}_{j,k}\oplus\mathbb{W}_{j,k},\quad\mathbb{V}_{j,k}\perp\mathbb{W}_{j,k}\end{gathered}$ (10)
@@
Assuming that each $\mathbb{V}_{J,k}$ is supported on $\mathcal{S}_{J,k}=\{v_{k}\}$ at the finest level $J$, we can define $\mathbb{V}_{J}:=\mathrm{span}\{\bm{e}_{k}\,|\,k\in[n]\}=\oplus_{k=1}^{n}\mathbb{V}_{J,k}$, where $\mathbb{V}_{J,k}:=\mathrm{span}(\{\bm{e}_{k}\})$ is simply the one-dimensional linear space associated with the vertex $v_{k}$ and generated by the $k$-th canonical basis vector $\bm{e}_{k}:=[0,\ldots,0,1,0\ldots,0]\in\mathbb{R}^{n}$. Note that by the parent-children relation, we can rewrite $\mathbb{V}_{J}$ as
$\mathbb{V}_{J}=\oplus_{k\in[n_{J-1}]}(\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}}).$ (11)
It is straightforward to see that (9) and (10) implies (7), which eventually implies (8) since we can decompose $\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}}$ as
$\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}}=\mathbb{V}_{J-1,k}\oplus\mathbb{W}_{J-1,k}$ (12)
for each $k\in[n_{J-1}]$, and the procedure can continue from bottom $j=J$ to top $j=0$. Hence, it remains to define a general procedure of decomposition for each non-leaf node in $\mathscr{T}_{J}({\mathcal{V}})$ such that (10) is satisfied. Our next result shows that we indeed can define such a general (bottom-up) procedure and obtain a tight framelet system for $\mathbb{V}_{J}$ based on Cororllary 1 and the partition tree $\mathscr{T}_{J}(\mathcal{V})$.
@@
Definition 2.
Let $\mathscr{T}_{J}({\mathcal{V}})$ be a partition tree as in Definition 1. Let $r_{j}$ be an integer such that $r_{j}\in[\min(|\mathcal{C}_{j,1}|,\dots,|\mathcal{C}_{j,n_{j}})|-1]$ and $\bm{A}^{[j,k]}\in\mathbb{R}^{r_{j}\times|\mathcal{C}_{j,k}|}$ and $\bm{B}^{[j,k]}\in\mathbb{R}^{m_{j,k}\times|\mathcal{C}_{j,k}|}$ be two filter matrices associated with the tree node $p_{j,k}$ for $j=0,\ldots,J-1$ and $k\in[n_{j}]$.
@@
    (Bottom-up Procedure) Recursively define at each level $j$ from $J-1$ to 00:
@@
      For each tree node $p_{j,k}$, obtain $\mathbb{V}_{j,k}:=\mathrm{span}(\bm{\Phi}^{[j,k]})$ and $\mathbb{W}_{j,k}=\mathrm{span}(\bm{\Psi}^{[j,k]})$ by Corollary 1 with subspace matrices $\{\bm{X}_{k^{\prime}}:=\bm{\Phi}^{[j+1,k^{\prime}]}\,|\,k^{\prime}\in\mathcal{C}_{j,k}\}$ and the two filter matrices $\bm{A}^{[j,k]}$ and $\bm{B}^{[j,k]}$. More precisely, let $\bm{X}:=\llbracket\bm{X}_{k^{\prime}}\rrbracket_{k^{\prime}\in\mathcal{C}_{j,k}}$. Then we have
      $\bm{\Phi}^{[j,k]}=\mathcal{A}^{[j,k]}\bm{X},\quad\bm{\Psi}^{[j,k]}=\mathcal{B}^{[j,k]}\bm{X},$ (13)
      where $\mathcal{A}^{[j,k]},\mathcal{B}^{[j,k]}$ are matrices corresponding to (6) in Corollary 1.
@@
    $\mathcal{F}_{J_{0}}^{J}(\mathcal{V}):=$ $\mathcal{F}_{J_{0}}(\{(\bm{A}^{[j,k]},\bm{B}^{[j,k]})\,|\,k\in[n_{j}]\}_{j=J_{0}}^{J-1})$ (14)
    ::=:= $\bm{\Phi}_{J_{0}}\cup\bm{\Psi}_{J_{0}}\cup\cdots\cup\bm{\Psi}_{J-1}.$
@@
Theorem 1.
Adopt the notations in Definition 2. Assume that the filter matrices $\bm{A}^{[j,k]}$ and $\bm{B}^{[j,k]}$ satisfy
$\bm{A}^{[j,k]}({\bm{A}^{[j,k]}})^{\top}$ $=\bm{I},$ (15)
$\bm{B}^{[j,k]}({\bm{A}^{[j,k]}})^{\top}$ $=\bm{0},$ (16)
$({\bm{B}^{[j,k]}})^{\top}\bm{B}^{[j,k]}$ $=\bm{I}-({\bm{A}^{[j,k]}})^{\top}\bm{A}^{[j,k]},$ (17)
@@
Proof.
We apply mathematical induction on $j$ (decreasingly) to prove that for each $j\in{0,\dots,J-1}$, all the subspaces $\mathbb{V}_{j+1,k^{\prime}}=\mathrm{span}(\bm{\Phi}^{[j+1,k^{\prime}]})$ in the previous level have the same dimension and conditions in Corrolary 1 are satisfied with $\bm{X}_{k}^{\prime}:=\bm{\Phi}^{[j+1,k^{\prime}]},k^{\prime}\in\mathcal{C}_{j,k}$. Hence, by the results of Corollary 1, we can obtain the conclusions of items (i)–(iii) of the theorem.
@@
    Induction hypothesis: Assume that at level $j+1$ with $10\leq j\leq J-$ 1, each $\mathbb{V}_{j+1,k}=\mathrm{span}(\bm{\Phi}^{[j+1,k]})$ has the same dimension $R_{j+1}:=\prod_{i=j+1}^{J}r_{i}$ for all $k\in[n_{j+1}]$, where each $\bm{\Phi}^{[j+1,k]}$ is formed by othonormal row vectors supported on $\mathcal{S}_{j+1,k}$.
@@
    Induction step: We next prove that each $\mathbb{V}_{j,k}=\mathrm{span}(\bm{\Phi}^{[j,k]})$ also has the same dimension $R_{j}:=\prod_{i=j}^{J}r_{i}$ for all $k\in[n_{j}]$ with each $\bm{\Phi}^{[j,k]}$ being formed by othonormal row vectors. In fact, fixed the node $p_{j,k}$, it has its children nodes $p_{j+1,k^{\prime}},k^{\prime}\in\mathcal{C}_{j,k}$. Consider $\bm{X}_{k^{\prime}}:=\bm{\Phi}^{[j+1,k^{\prime}]}\in\mathbb{R}^{R_{j+1}\times n}$, $k^{\prime}\in\mathcal{C}_{j,k}$. By (15) – (17), the conditions of Corollary 1 are satisfied with these subspace matrices $\bm{X}_{k^{\prime}}$ and the two filter matrices $\bm{A}^{[j,k]},\bm{B}^{[j,k]}$. Thus its conclusions can be obtained. That is, $\mathbb{V}_{\bm{X}}:=\oplus_{k^{\prime}\in\mathcal{C}_{j,k}}\mathbb{V}_{j+1,k^{\prime}}=\mathbb{V}_{j,k}\oplus\mathbb{W}_{j,k},\mathbb{V}_{j,k}\perp\mathbb{W}_{j,k}$, and the rows of $\bm{\Phi}^{[j,k]}$ form an orthonormal basis for $\mathbb{V}_{j,k}$. Moreover, the rows of $\bm{\Psi}^{[j,k]}$ form a tight frame for $\mathbb{W}_{j,k}$. Note that the dimension of $\mathbb{V}_{j,k}$ satisfies $\mathrm{dim}(\mathbb{V}_{j,k})\equiv r_{j}R_{j+1}=\prod_{i=j}^{J}r_{i}$ for all $k\in[n_{j}]$.
@@
Consequently, item (i) holds and it implies item (ii). Item (iii) follows from item (ii) and the result that the rows of $\bm{\Psi}^{[j,k]}$ form a tight frame for $\mathbb{W}_{j,k}$. ∎
The construction of $\mathcal{V}$-framelet system in Theorem 1 is formulated in Algorithm 2. Before we proceed, we have the following remarks.
@@
where $m$ is uniquely determined by the integers $s,t$ with $1\leq s<t\leq|\mathcal{C}_{j,k}|$ as $m:=m(s,t,|\mathcal{C}_{j,k}|):=\frac{(2|\mathcal{C}_{j,k}|-s)(s-1)}{2}+s-t$. It is shown that (17) in Theorem 1 is satisfied. In summary, the rows of $\bm{B}^{[j,k]}$ is constructed by enumerating pairs of indices in $\mathcal{C}_{j,k}$ and assigning the same constant with different signs, which resemble classic Haar wavelets/filters. Our general construction includes [28] as a special case.
II-C Construction of $\mathcal{G}$-Framelet Systems
As we point out in Remark 2, the $\mathcal{V}$-framelet systems do not involve the other graph structure such as the edge set $\mathcal{E}$. However, they do depend on the partition tree $\mathscr{T}_{J}(\mathcal{V})$ and how to obtain such a partition tree is closely related to the graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$. We next discuss how we can obtain the partition tree from a given graph $\mathcal{G}$ and define the so-called graph-involved $\mathcal{G}$-framelet systems.

### v3

@@ 2.1 Preliminaries
(2) $f=\sum_{i\in\mathcal{I}}\langle f,\phi_{i}\rangle_{\mathcal{H}}\phi_{i},\quad\forall f\in\mathcal{H}.$
@@
Proof 2.2.
[…]
Lemma 2.1 can be extended to the following more general result, which will be useful in our later construction of graph framelet systems.
Corollary 2.3.
@@
(3) $\mathbb{V}_{\bm{X}}$ $:=\mathrm{span}(\bm{X})\,\,with\,\,\bm{X}:=\llbracket\bm{X}_{i}\rrbracket_{i\in[m]},$
(4) $\mathbb{V}$ $:=\mathrm{span}(\bm{\Phi})\,\,\,with\,\,\bm{\Phi}:=\llbracket\bm{\Phi}_{\ell}\rrbracket_{\ell\in[r]},$
(5) $\mathbb{W}$ $:=\mathrm{span}(\bm{\Psi})\,\,\,with\,\,\bm{\Psi}:=\llbracket\bm{\Psi}_{\ell}\rrbracket_{\ell\in[r]}.$
Proof 2.4.
@@
(7) $\left\{\begin{matrix}\mathcal{A}:=\mathrm{diag}(\bm{A},\ldots,\bm{A})\bm{E}\in\mathbb{R}^{m_{1}r\times mr},\\ \mathcal{B}:=\mathrm{diag}(\bm{B},\ldots,\bm{B})\bm{E}\in\mathbb{R}^{m_{2}r\times mr}.\end{matrix}\right$.
Then, we have $\mathcal{A}\mathcal{A}^{\top}=\bm{I}$, $\mathcal{B}\mathcal{A}^{\top}=\bm{0}$, and $\mathcal{B}^{\top}\mathcal{B}=\bm{I}-\mathcal{A}^{\top}\mathcal{A}$. Moreover, $\bm{\Phi}=\mathcal{A}\bm{X}$ and $\bm{\Psi}=\mathcal{B}\bm{X}$. The conclusion then directly follows from applying Lemma 2.1.
Remark 2.5.
We call the matrices $\bm{A},\bm{B}$ in Corollary 2.3 as filters or filter matrices, and the matrices $\bm{X}_{i},i\in[m]$ as subspace matrices. We formulate Corollary 2.3 as an algorithm in Algorithm 1. We highlight that the key of Corollary 2.3 is that each subspace $\mathrm{span}(\bm{X}_{i})$ has the same dimension $r$ for all $i\in[m]$. Such a condition will play a crucial role in our later construction of tight graph framelet systems.
@@
[…]
@@
2.2 Construction of $\mathcal{V}$-Framelet Systems
@@
(8) $\mathbb{V}_{j}=\mathbb{V}_{j-1}\oplus\mathbb{W}_{j-1},\quad\mathbb{V}_{j-1}\perp\mathbb{W}_{j-1},\quad j\in[J],$
then by doing the decomposition J1J-1 times, we have
(9) $\mathbb{V}_{J}=\mathbb{V}_{0}\oplus\mathbb{W}_{0}\oplus\mathbb{W}_{1}\oplus\dots\oplus\mathbb{W}_{J-1}$
@@
(10) $\mathbb{V}_{j}:=\oplus_{k=1}^{n_{j}}\mathbb{V}_{j,k},\quad\mathbb{W}_{j}:=\oplus_{k=1}^{n_{j}}\mathbb{W}_{j,k},$
@@
(11) $\begin{gathered}\oplus_{k^{\prime}\in\mathcal{C}_{j,k}}\mathbb{V}_{j+1,k^{\prime}}=\mathbb{V}_{j,k}\oplus\mathbb{W}_{j,k},\quad\mathbb{V}_{j,k}\perp\mathbb{W}_{j,k}\end{gathered}$
@@
Assuming that each $\mathbb{V}_{J,k}$ is supported on $\mathcal{S}_{J,k}=\{v_{k}\}$ at the finest level $J$, we can define $\mathbb{V}_{J}:=\mathrm{span}\{\bm{e}_{k}\,|\,k\in[n]\}=\oplus_{k=1}^{n}\mathbb{V}_{J,k}$, where $\mathbb{V}_{J,k}:=\mathrm{span}(\{\bm{e}_{k}\})$ is simply the one-dimensional linear space associated with the vertex $v_{k}$ and generated by the $k$-th canonical basis vector $\bm{e}_{k}:=[0,\ldots,0,1,0\ldots,0]\in\mathbb{R}^{n}$. Note that by the parent-children relation, we can rewrite $\mathbb{V}_{J}$ as $\mathbb{V}_{J}=\oplus_{k\in[n_{J-1}]}(\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}})$. It is straightforward to see that (10) and (11) implies (8), which eventually implies (9) since we can decompose $\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}}$ as $\oplus_{k^{\prime}\in\mathcal{C}_{J-1,k}}\mathbb{V}_{J,k^{\prime}}=\mathbb{V}_{J-1,k}\oplus\mathbb{W}_{J-1,k}$ for each $k\in[n_{J-1}]$, and the procedure can continue from bottom $j=J$ to top $j=0$. Hence, it remains to define a general procedure of decomposition for each non-leaf node in $\mathscr{T}_{J}({\mathcal{V}})$ such that (11) is satisfied. Our next result shows that we indeed can define such a general (bottom-up) procedure and obtain a tight framelet system for $\mathbb{V}_{J}$ based on Cororllary 2.3 and the partition tree $\mathscr{T}_{J}(\mathcal{V})$.
@@
Definition 2.7.
Let $\mathscr{T}_{J}({\mathcal{V}})$ be a partition tree as in Definition 2.6. Let $r_{j}$ be an integer such that $r_{j}\in[\min(|\mathcal{C}_{j,1}|,\dots,|\mathcal{C}_{j,n_{j}})|-1]$ and $\bm{A}^{[j,k]}\in\mathbb{R}^{r_{j}\times|\mathcal{C}_{j,k}|}$ and $\bm{B}^{[j,k]}\in\mathbb{R}^{m_{j,k}\times|\mathcal{C}_{j,k}|}$ be two filter matrices associated with the tree node $p_{j,k}$ for $j=0,\ldots,J-1$ and $k\in[n_{j}]$.
@@
    (Bottom-up Procedure) Recursively define at each level $j$ from J1J-1 to 0:
@@
      For each tree node $p_{j,k}$, obtain $\mathbb{V}_{j,k}:=\mathrm{span}(\bm{\Phi}^{[j,k]})$ and $\mathbb{W}_{j,k}=\mathrm{span}(\bm{\Psi}^{[j,k]})$ by Corollary 2.3 with subspace matrices $\{\bm{X}_{k^{\prime}}:=\bm{\Phi}^{[j+1,k^{\prime}]}\,|\,k^{\prime}\in\mathcal{C}_{j,k}\}$ and the two filter matrices $\bm{A}^{[j,k]}$ and $\bm{B}^{[j,k]}$. More precisely, let $\bm{X}:=\llbracket\bm{X}_{k^{\prime}}\rrbracket_{k^{\prime}\in\mathcal{C}_{j,k}}$. Then we have
      (12) $\bm{\Phi}^{[j,k]}=\mathcal{A}^{[j,k]}\bm{X},\quad\bm{\Psi}^{[j,k]}=\mathcal{B}^{[j,k]}\bm{X},$
      where $\mathcal{A}^{[j,k]},\mathcal{B}^{[j,k]}$ are matrices corresponding to (7) in Corollary 2.3.
@@
    (13) $\mathcal{F}_{J_{0}}^{J}(\mathcal{V})$= $\mathcal{F}_{J_{0}}(\{(\bm{A}^{[j,k]},\bm{B}^{[j,k]})\,|\,k\in[n_{j}]\}_{j=J_{0}}^{J-1})$
    = $\bm{\Phi}_{J_{0}}\cup\bm{\Psi}_{J_{0}}\cup\cdots\cup\bm{\Psi}_{J-1}.$
@@
Theorem 2.8.
Adopt the notations in Definition 2.7. Assume that the filter matrices $\bm{A}^{[j,k]}$ and $\bm{B}^{[j,k]}$ satisfy
(14) $\bm{A}^{[j,k]}({\bm{A}^{[j,k]}})^{\top}$ $=\bm{I},$
(15) $\bm{B}^{[j,k]}({\bm{A}^{[j,k]}})^{\top}$ $=\bm{0},$
(16) $({\bm{B}^{[j,k]}})^{\top}\bm{B}^{[j,k]}$ $=\bm{I}-({\bm{A}^{[j,k]}})^{\top}\bm{A}^{[j,k]},$
@@
Proof 2.9.
We apply mathematical induction on $j$ (decreasingly) to prove that for each $j\in{0,\dots,J-1}$, all the subspaces $\mathbb{V}_{j+1,k^{\prime}}=\mathrm{span}(\bm{\Phi}^{[j+1,k^{\prime}]})$ in the previous level have the same dimension and conditions in Corrolary 2.3 are satisfied with $\bm{X}_{k}^{\prime}:=\bm{\Phi}^{[j+1,k^{\prime}]},k^{\prime}\in\mathcal{C}_{j,k}$. Hence, by the results of Corollary 2.3, we can obtain the conclusions of items (i)–(iii) of the theorem.
@@
    Induction hypothesis: Assume that at level $j+1$ with $0\leq j\leq J-1$, each $\mathbb{V}_{j+1,k}=\mathrm{span}(\bm{\Phi}^{[j+1,k]})$ has the same dimension $R_{j+1}:=\prod_{i=j+1}^{J}r_{i}$ for all $k\in[n_{j+1}]$, where each $\bm{\Phi}^{[j+1,k]}$ is formed by othonormal row vectors supported on $\mathcal{S}_{j+1,k}$.
@@
    Induction step: We next prove that each $\mathbb{V}_{j,k}=\mathrm{span}(\bm{\Phi}^{[j,k]})$ also has the same dimension $R_{j}:=\prod_{i=j}^{J}r_{i}$ for all $k\in[n_{j}]$ with each $\bm{\Phi}^{[j,k]}$ being formed by othonormal row vectors. In fact, fixed the node $p_{j,k}$, it has its children nodes $p_{j+1,k^{\prime}},k^{\prime}\in\mathcal{C}_{j,k}$. Consider $\bm{X}_{k^{\prime}}:=\bm{\Phi}^{[j+1,k^{\prime}]}\in\mathbb{R}^{R_{j+1}\times n}$, $k^{\prime}\in\mathcal{C}_{j,k}$. By (14) – (16), the conditions of Corollary 2.3 are satisfied with these subspace matrices $\bm{X}_{k^{\prime}}$ and the two filter matrices $\bm{A}^{[j,k]},\bm{B}^{[j,k]}$. Thus its conclusions can be obtained. That is, $\mathbb{V}_{\bm{X}}:=\oplus_{k^{\prime}\in\mathcal{C}_{j,k}}\mathbb{V}_{j+1,k^{\prime}}=\mathbb{V}_{j,k}\oplus\mathbb{W}_{j,k},\mathbb{V}_{j,k}\perp\mathbb{W}_{j,k}$, and the rows of $\bm{\Phi}^{[j,k]}$ form an orthonormal basis for $\mathbb{V}_{j,k}$. Moreover, the rows of $\bm{\Psi}^{[j,k]}$ form a tight frame for $\mathbb{W}_{j,k}$. Note that the dimension of $\mathbb{V}_{j,k}$ satisfies $\mathrm{dim}(\mathbb{V}_{j,k})\equiv r_{j}R_{j+1}=\prod_{i=j}^{J}r_{i}$ for all $k\in[n_{j}]$.
@@
Consequently, item (i) holds and it implies item (ii). Item (iii) follows from item (ii) and the result that the rows of $\bm{\Psi}^{[j,k]}$ form a tight frame for $\mathbb{W}_{j,k}$.
The construction of $\mathcal{V}$-framelet system in Theorem 2.8 is formulated in Algorithm 2. Before we proceed, we have the following remarks.
@@
where $m$ is uniquely determined by the integers $s,t$ with $1\leq s<t\leq|\mathcal{C}_{j,k}|$ as $m:=m(s,t,|\mathcal{C}_{j,k}|):=\frac{(2|\mathcal{C}_{j,k}|-s)(s-1)}{2}+s-t$. It is shown that (16) in Theorem 2.8 is satisfied. In summary, the rows of $\bm{B}^{[j,k]}$ is constructed by enumerating pairs of indices in $\mathcal{C}_{j,k}$ and assigning the same constant with different signs, which resemble classic Haar wavelets/filters. Our general construction includes [35] as a special case.
2.3 Construction of $\mathcal{G}$-Framelet Systems
As we point out in Remark 2.10, the $\mathcal{V}$-framelet systems do not involve the other graph structure such as the edge set $\mathcal{E}$. However, they do depend on the partition tree $\mathscr{T}_{J}(\mathcal{V})$ and how to obtain such a partition tree is closely related to the graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$. We next discuss how we can obtain the partition tree from a given graph $\mathcal{G}$ and define the so-called graph-involved $\mathcal{G}$-framelet systems.